Recently, the skeleton of King Richard III was found under a parking lot in England. If tissue samples from the skeleton, analyzed in 2014, contain about 93.79% of the carbon-14 expected in living tissue, what year did King Richard III die? The half-life for carbon-14 is 5715 years.
Answer #1

Given:

Half life = 5.715*10^3 years

use relation between rate constant and half life of 1st order reaction

k = (ln 2) / k

= 0.693/(half life)

= 0.693/(5.715*10^3)

= 1.213*10^-4 years-1

we have:

[A]o = 100

[A] = 93.79

k = 1.213*10^-4 years-1

use integrated rate law for 1st order reaction

ln[A] = ln[A]o - k*t

ln(93.79) = ln(100) - 1.213*10^-4*t

4.541 = 4.605 - 1.213*10^-4*t

1.213*10^-4*t = 6.411*10^-2

t = 5.287*10^2 years

Answer: 529 years